# End User Authoring of Personalized Content Classifiers: Comparing Example Labeling, Rule Writing, and LLM Prompting

### v1

@@ Abstract.
Existing tools for laypeople to create personal classifiers often assume a motivated user working uninterrupted in a single, lengthy session. However, users tend to engage with social media casually, with many short sessions on an ongoing, daily basis. To make creating personal classifiers for content curation easier for such users, tools should support rapid initialization and iterative refinement. In this work, we compare three strategies—(1) example labeling, (2) rule writing, and (3) large language model (LLM) prompting—for end users to build personal content classifiers. From an experiment with 37 non-programmers tasked with creating personalized comment moderation filters, we found that with LLM prompting, participants reached 95% of peak performance in 5 minutes, beating other strategies due to higher recall, but all strategies struggled with iterative refinement. Despite LLM prompting’s better performance, participants preferred different strategies in different contexts and, even when prompting, provided examples or wrote rule-like prompts, suggesting hybrid approaches.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: The ACM CHI conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, Hawai’iprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Collaborative and social computing systems and tools
@@ 1. Introduction
In this work, we use these system requirements to compare three prominent strategies for creating custom classifiers for content curation: (1) labeling examples for supervised learning, (2) writing and carrying out rules, and (3) prompting a large language model (LLM). Each of these techniques has its respective benefits and drawbacks for end-user classifier creation. In interactive machine learning (IML) research, labeling examples is a popular strategy to customize a classifier. It is promising in the social media context due to its simplicity and ability to be easily divided into smaller tasks. However, even though pretrained word-embeddings (Devlin et al., 2018) and active learning (Settles, 2009) reduce the number of labels needed to train a classifier, labeling examples can still be tedious and inefficient (Das et al., 2013; Kulesza et al., 2014), so this technique is typically not available for social media users to create personal classifiers from scratch. In contrast, the most common strategy deployed on social platforms today is writing rules. Most platforms allow users to curate content through simple keywords (Jhaver et al., 2022). AutoMod, the most widely used content moderation tool on Reddit (Jhaver et al., 2019; Chandrasekharan et al., 2019a), further enables community moderators to write regular expressions to remove inappropriate posts (Jhaver et al., 2019). Despite the transparency that rules afford, users often struggle to write rules that account for nuanced contexts and to further refine them (Jhaver et al., 2022; Song et al., 2023; Chandrasekharan et al., 2019a). Finally, with recent advances in LLMs, we additionally investigate the technique of customizing zero-shot or few-shot classifiers by prompting an LLM in natural language (Kumar et al., 2023; OpenAI, 2024). While LLMs have the potential to learn nuanced preferences quickly (Wang et al., 2019), their effective use may require prompt engineering skills beyond the capabilities of many end users (Sclar et al., 2023; Lu et al., 2021).
@@
We found that writing prompts generally allowed participants to create personal classifiers with higher performance more quickly, achieving a significantly higher recall and $F_{1}$ score than authoring rules and labeling examples throughout the 15-minute experiment period. However, participants struggled to iteratively refine their prompts to align LLMs with their nuanced preferences, as both recall and precision plateaued five minutes or earlier after starting. This difficulty was most pronounced when users tried to teach LLMs to categorize specific phrases within certain concepts, such as classifying “fool” as offensive or “hell” as obscene. Even worse, the opaque nature and occasionally unpredictable behavior of LLMs left participants uncertain about which phrases LLMs might interpret differently. As a result, one-third of participants preferred labeling examples over writing prompts, while another third preferred authoring rules. Some found labeling examples to be the easiest method for expressing their intuitive preferences, while they considered rules better for curating content about specific topics or events. Participants even tried to deploy these two strategies when iterating on their prompts: they often added incorrectly classified examples directly into their prompts as few-shot examples or wrote prompts resembling rules.
Our findings offer a roadmap for future end-user content curation systems and have broader implications for facilitating more efficient and effective communication between end users and classifiers. First, LLMs are a promising tool for enabling personalized content classifiers, with only five minutes of work needed to initialize a classifier with a better $F_{1}$ score than 15 minutes using traditional methods of labeling examples or writing rules. However, we find that labeling examples requires less cognitive effort from users, and writing rules provides users with more control and transparency over their classifiers. Given the diversity of content curation scenarios and user preferences, future content curation systems should provide end users with more flexible strategies to create and iterate on their classifiers. For instance, while writing prompts could be the default strategy given its superior overall performance, future hybrid systems could also suggest possible prompts based on example labels to help users who have difficulty articulating their intuitions, or allow rule authoring for preferences about specific topics or events. Hybrid approaches such as these could bring together the benefits of all three strategies and facilitate more effective human-LLM collaboration.
@@ 2.1. End User Customization of Content Curation
Content curation is an important process for internet users who seek to classify desired or undesired content on social media (Lee et al., 2019; Jhaver et al., 2023; Klug et al., 2023), filter out harassment or spam from their email folders (Park et al., 2019; Mahar et al., 2018), manage online chats (Sophia C. Volk and Strauß, 2024; Wang et al., 2023b), find relevant news articles (Sophia C. Volk and Strauß, 2024; Josephine B. Schmitt and Schneider, 2018), or otherwise categorize information that they encounter online. However, content curation tools are often centralized and platform-wide, leaving limited room for end user customization (Neves and Reis, 2018). For example, one-size-fits-all moderation algorithms often fail to account for the diverse perceptions of toxicity across countries (Jiang et al., 2021) and communities (Weld et al., 2022). Similarly, some users want to categorize their emails into more fine-grained folders other than the default “Spam” or “Promotions” folders (Park et al., 2019; Yoo et al., 2011). As a result, there have been growing calls to empower end users to customize content curation classifiers. For example, researchers have argued for personal content moderation tools that enable users to customize some aspects of their moderation preferences based on the content of posts uploaded by other users (Jhaver et al., 2023; Jhaver and Zhang, 2023). Users may wish to remove unwanted content from their feeds on platforms like Instagram, TikTok, and X/Twitter (Jhaver et al., 2023) or to moderate the comment sections of their YouTube videos or the chat channels of their Twitch streams (Jhaver et al., 2022). In both cases, end users act as moderators that remove content according to their preferences.
@@
Researchers and practitioners have experimented with offering custom content curation tools. However, these tools often struggle to navigate the trade-off between flexibility of customization and ease of use. Some tools offer extensive customization but are too complex for casual users. For example, Reddit’s AutoMod (Jhaver et al., 2019) allows community moderators to write regular expressions in YAML to remove inappropriate posts, but only users with technical knowledge of regular expressions and programming syntax can fully utilize AutoMod’s flexibility (Song et al., 2023). Conversely, other tools are more simplistic in design. Many of them involve pre-trained classifiers for platform-defined or researcher-defined concepts such as racism, misogyny, and political views, with users only able to adjust their sensitivity to each concept (Chandrasekharan et al., 2019a; Instagram, 2024; Bhargava et al., 2019; Bodyguard, 2024). However, this approach fails when users have differing definitions of a concept or wish to classify content with respect to a concept that is not provided (Jhaver et al., 2023). Recently, research has shown that LLMs can outperform state-of-the-art classifiers in detecting toxic content with natural language prompts (Kumar et al., 2023; OpenAI, 2024). But LLMs are not exempt from the customization versus ease of use trade-off: when applied toward more complex and nuanced curation preferences that go beyond simple hate speech or toxicity filtering, LLMs sometimes perform even worse than a coin toss (Kumar et al., 2023). In this work, we compare common content curation techniques in terms of this trade-off through an experiment using the case of personal content moderation, with the eventual goal of designing a system that offers both customization and ease of use for a broad range of end users.
@@ 2.2. Strategies that Support End Users in Customizing Classifiers for Content Curation
Despite the ubiquitous use of machine learning (ML) algorithms in various domains, only a small group of people with technical expertise possess the skills to develop these algorithms. Interactive machine learning (IML) seeks to democratize ML training with humans-in-the-loop, enabling non-experts to participate through “rapid, focused, and incremental model updates” (Amershi et al., 2014). IML underlies many efforts at developing end user systems to build classifiers and operates in three critical phases  (Amershi, 2011; Stumpf et al., 2009): (1) Users examine a model’s outputs across a set of examples; (2) Users then offer feedback to the model, guiding its learning in the desired direction; and (3) Users assess the model’s overall performance and decide whether to conclude the training process or to offer the model more feedback.
At the heart of these phases are teaching vocabularies: the frameworks through which end users structure their feedback  (Porter et al., 2013), including assigning labels, selecting features, specifying parameters, or indicating error preferences (Kapoor et al., 2010). Recently, prompts have emerged as another promising way for end users to communicate with algorithms (i.e., LLM-based models). Through the lens of teaching vocabularies, we characterize existing content curation systems by three primary strategies that they leverage to support end user customization: labels, features or rules, and prompts. In the following, we discuss each strategy and its relationship to content curation.
@@ 2.2.2. Features or Rules
In response to the aforementioned criticisms of the labeling approach, a body of research focuses on feature-level human input (Hema et al., 2006; Wallace et al., 2012; Šavelka et al., 2015; Cheng and Bernstein, 2015). Here, the term “feature” refers to an attribute of a data instance that ML algorithms use to predict its class label. In comparison, “rules” are often logical statements that directly map to a class label without relying on ML algorithms. In systems that support technical users to build a text classifier via feature-level input, features and rules often take the same form: the mentions of keywords. By incorporating features or rules into different algorithm architectures, these systems can offer varying degrees of transparency. Rules can offer the highest transparency by directly classifying text without using any ML models. Features, on the other hand, can be incorporated into transparent ML algorithms, such as Decision Trees, Naive Bayes algorithms (Kulesza et al., 2015), and Support Vector Machines (Šavelka et al., 2015). In addition, because end users might produce noisy and inconsistent features (Das et al., 2013; Boecking et al., 2020), user-generated features can also be used to label data. Weak supervision algorithms, for example, then learn from these “soft labels” (Ratner et al., 2017). However, these indirect approaches sacrifice the transparency and control that users often assume with feature-level input. For this reason, many systems designed for end users often integrate features with transparent algorithms to enhance users’ mental models of such systems (Kulesza et al., 2015; Šavelka et al., 2015).
Although empirical results regarding whether feature-level input from end users improves performance have been mixed (Hema et al., 2006; Das et al., 2013; Wu et al., 2019), researchers are generally in agreement that feature-level input requires fewer user actions to achieve comparable results to example labeling, and that it could produce models more closely aligned with an individual’s needs or domain knowledge. However, features can be seen as too granular and therefore not generalizable enough, posing a challenge for end users in creating effective classifiers (Chandrasekharan et al., 2019b; Song et al., 2023).
@@ 2.2.3. Prompts
LLMs have exploded in popularity due to their proficiency in a broad range of natural language processing tasks, such as sentiment analysis and machine translation. Compared to features, natural language prompts can convey richer user guidance to models and thus promote greater generalizability across various contexts (Sahoo et al., 2024). Rather than having to train a new model for every custom task, users can simply customize LLMs by feeding them prompts at run time. Such an ability to recognize the desired task on-the-fly is called in-context learning (ICL) (Brown et al., 2020).
Developing effective prompts for ICL is crucial to leveraging LLMs in content curation (Min et al., 2022; Mishra et al., 2021). To date, the most common patterns for prompting are zero-shot or few-shot prompts. Zero-shot prompts give instructions for a task without any specific examples for training (Wang et al., 2019). Research indicates that the performance of zero-shot prompts can be enhanced by iteratively refining task instructions (Efrat and Levy, 2020) or breaking down tasks into simpler subtasks (Wei et al., 2022). Alternatively, few-shot prompts incorporate a few input-output examples to showcase the desired pattern to which LLMs should adhere (Sahoo et al., 2024; Brown et al., 2020). The quality of few-shot prompts relies heavily on the selected examples  (Rubin et al., 2021). While few-shot prompts generally outperform zero-shot prompts, the simplicity of composing natural language instructions without the need for selecting examples makes zero-shot prompting an attractive option for end users. Even so, various challenges impede end users in creating and refining their prompts. For example, even a slight adjustment to a prompt’s format or ordering of examples can greatly influence a model’s performance (Sclar et al., 2023; Lu et al., 2021). In addition, users often find it difficult to evaluate the clarity of their instructions to LLMs and to align their perspectives with those of LLMs, leading to incorrect or unexpected outputs (Deng et al., 2023).
@@ 2.3. How Existing IML Systems Overlook End User Needs in Content Curation
While researchers and practitioners have built many systems using labels, features, and prompts, they often make assumptions about end users that do not translate well into the context of content curation.

### v2

@@ Abstract.
Existing tools for laypeople to create personal classifiers often assume a motivated user working uninterrupted in a single, lengthy session. However, users tend to engage with social media casually, with many short sessions on an ongoing, daily basis. To make creating personal classifiers for content curation easier for such users, tools should support rapid initialization and iterative refinement. In this work, we compare three strategies—(1) example labeling, (2) rule writing, and (3) large language model (LLM) prompting—for end users to build personal content classifiers. From an experiment with 37 non-programmers tasked with creating personalized moderation filters, we found that participants preferred different initializing strategies in different contexts, despite LLM prompting’s better performance. However, all strategies faced challenges with iterative refinement. To overcome challenges in iterating on their prompts, participants even adopted hybrid approaches such as providing examples as in-context examples or writing rule-like prompts.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: The ACM CHI conference on Human Factors in Computing Systems; April 26–May 1st, 2024; Yokohama, Japanprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Collaborative and social computing systems and tools
@@ 1. Introduction
In this work, we use these system requirements to compare three prominent strategies for creating custom classifiers for content curation: (1) labeling examples for supervised learning, (2) writing and carrying out rules, and (3) prompting a large language model (LLM). In particular, we aim to answer the following two research questions:
  • Which strategy allows end users to initialize a personal content classifier the most easily?
  • Which strategy allows end users to iterate on their personal content classifiers after initialization the most easily?
Each of these three strategies has its respective benefits and drawbacks for end-user classifier creation. In interactive machine learning (IML) research, labeling examples is a popular strategy to customize a classifier. It is promising in the social media context due to its simplicity and ability to be easily divided into smaller tasks. However, even though pretrained word-embeddings (Devlin et al., 2018) and active learning (Settles, 2009) reduce the number of labels needed to train a classifier, labeling examples can still be tedious and inefficient (Das et al., 2013; Kulesza et al., 2014), so this technique is typically not available for social media users to create personal classifiers from scratch. In contrast, the most common strategy deployed on social platforms today is writing rules. Most platforms allow users to curate content through simple keywords (Jhaver et al., 2022). AutoMod, the most widely used content moderation tool on Reddit (Jhaver et al., 2019; Chandrasekharan et al., 2019a), further enables community moderators to write regular expressions to remove inappropriate posts (Jhaver et al., 2019). Despite the transparency that rules afford, users often struggle to write rules that account for nuanced contexts and to further refine them (Jhaver et al., 2022; Song et al., 2023; Chandrasekharan et al., 2019a). Finally, with recent advances in LLMs, we additionally investigate the technique of customizing zero-shot or few-shot classifiers by prompting an LLM in natural language (Kumar et al., 2023; OpenAI, 2024). While LLMs have the potential to learn nuanced preferences quickly (Wang et al., 2019), their effective use may require prompt engineering skills beyond the capabilities of many end users (Sclar et al., 2023; Lu et al., 2021).
@@
Through a quantitative and qualitative analysis, we compared how these three strategies support end users to initialize and iterate on their personal content filters. In more detail, we present the following findings:
  • We found that writing prompts generally allowed participants to create personal content filters with higher performance more quickly, primarily driven by having the highest recall despite comparable precision across conditions.
  • Nonetheless, participants expressed reason to prefer each of the strategies when initializing a filter:
    • When preferences were ill-defined but intuitive, labeling examples was considered the easiest method;
    • When participants had specific topics or events they were trying to filter, writing rules was considered easiest; and
    • When preferences were both relatively well-defined and general, writing prompts was considered easiest.
  • All strategies struggled with iterative refinement. Participants particularly found it challenging aligning LLMs with their nuanced preferences, as recall and precision often plateaued within the first five minutes.
  • In their effort to improve their filters while writing prompts, participants often turned to approaches that looked like the other strategies, for instance, by adding misclassified examples directly as in-context examples or writing prompts resembling rules.
  • Finally, no single strategy was clearly preferred across the board. We catalog participants’ needs beyond ease that led to a divergence in preferences, including their varying tolerance for viewing toxic content and prioritization of precision versus recall.
Our findings offer a roadmap for future end-user content curation systems and have broader implications for facilitating more efficient and effective communication between end users and classifiers. First, while LLMs are a promising tool for enabling personalized content classifiers, we find that labeling examples requires less cognitive effort from users, and writing rules provides users with more control and transparency over their classifiers. Given the diversity of content curation scenarios and user preferences, future content curation systems should provide end users with more flexible strategies to create and iterate on their classifiers. For instance, while writing prompts could be the default strategy given its superior overall performance, future hybrid systems could also suggest possible prompts based on example labels to help users who have difficulty articulating their intuitions, or allow rule authoring for preferences about specific topics or events. Hybrid approaches such as these could bring together the benefits of all three strategies and facilitate more effective human-LLM collaboration.
In summary, we made the following two contributions. First, this paper explores how existing interactive machine learning strategies fail to address the diverse and dynamic user needs of content curation through a rigorous within-subject experiment. Second, it provides a nuanced perspective on the use of LLMs in content moderation. This paper highlights their limitations in bootstrapping and iteration and offers design implications to better support effective human-LLM collaboration.
@@ 2.1. End User Customization of Content Curation
Content curation is an important process for internet users who seek to classify desired or undesired content on social media (Lee et al., 2019; Jhaver et al., 2023; Klug et al., 2023), filter out harassment or spam from their email folders (Park et al., 2019; Mahar et al., 2018), manage online chats (Sophia C. Volk and Strauß, 2024; Wang et al., 2023b), or find relevant news articles (Sophia C. Volk and Strauß, 2024; Josephine B. Schmitt and Schneider, 2018). However, content curation tools are often centralized and platform-wide, leaving limited room for end-user customization (Neves and Reis, 2018). For example, one-size-fits-all moderation algorithms often fail to account for the diverse perceptions of toxicity across countries (Jiang et al., 2021) and communities (Weld et al., 2022). Similarly, some users want to categorize their emails into more fine-grained folders other than the default “Spam” or “Promotions” folders (Park et al., 2019; Yoo et al., 2011). As a result, there have been growing calls to empower end users to customize content curation classifiers (Jhaver et al., 2023; Jhaver and Zhang, 2023). In particular, researchers have argued for personal content moderation tools that enable users to customize their social media feeds. Users may wish to remove unwanted content from their feeds on platforms like Instagram, TikTok, and X/Twitter (Jhaver et al., 2023) or to moderate the comment sections of their YouTube videos or their Twitch streams (Jhaver et al., 2022). In both cases, end users act as moderators that remove content according to their preferences.
@@
Researchers and practitioners have experimented with offering custom content curation tools. However, these tools often struggle to navigate the trade-off between flexibility of customization and ease of use. Some tools offer extensive customization but are too complex for casual users. For example, Reddit’s AutoMod (Jhaver et al., 2019) allows community moderators to write regular expressions in YAML to remove inappropriate posts, but only users with technical knowledge of regular expressions and programming syntax can fully utilize AutoMod’s flexibility (Song et al., 2023). Conversely, other tools are more simplistic in design. Many of them involve pre-trained classifiers for platform-defined or researcher-defined concepts such as racism, misogyny, and political views, with users only able to adjust their sensitivity to each concept (Chandrasekharan et al., 2019a; Instagram, 2024; Bhargava et al., 2019; Bodyguard, 2024). However, this approach fails when users have differing definitions of a concept or have a new concept (Jhaver et al., 2023). Recently, research has shown that LLMs can outperform state-of-the-art classifiers in detecting toxic content with natural language prompts (Kumar et al., 2023; OpenAI, 2024). But LLMs are not exempt from the customization versus ease of use trade-off: when applied toward more complex and nuanced curation preferences that go beyond simple hate speech or toxicity filtering, LLMs sometimes perform even worse than a coin toss (Kumar et al., 2023). In this work, we compare common content curation techniques in terms of this trade-off through an experiment using the case of personal content moderation, with the eventual goal of designing a system that offers both customization and ease of use for a broad range of end users.
@@ 2.2. Strategies that Support End Users in Customizing Classifiers for Content Curation
Interactive machine learning (IML) seeks to democratize ML training with humans-in-the-loop, enabling non-experts to participate through “rapid, focused, and incremental model updates” (Amershi et al., 2014). IML enables end users to build classifiers in three critical phases  (Amershi, 2011; Stumpf et al., 2009): (1) Users examine a model’s outputs across a set of examples; (2) Users then offer feedback to the model, guiding its learning in the desired direction; and (3) Users assess the model’s overall performance and decide whether to conclude the training process or to offer the model more feedback.
At the heart of these phases are teaching vocabularies: the frameworks through which end users structure their feedback (Porter et al., 2013), including assigning labels, selecting features, or indicating error preferences (Kapoor et al., 2010). Recently, prompts have emerged as another promising way for end users to communicate with LLM-based models. Through the lens of teaching vocabularies, we characterize existing content curation systems by three primary strategies that they leverage to support end-user customization: labels, features or rules, and prompts. In the following, we discuss each strategy and its relationship to content curation.
@@ 2.2.2. Features or Rules
In response to the aforementioned criticisms of the labeling approach, a body of research focuses on feature-level human input (Hema et al., 2006; Wallace et al., 2012; Šavelka et al., 2015; Cheng and Bernstein, 2015). Here, the term “feature” refers to an attribute of a data instance that ML algorithms use to predict its class label. In comparison, “rules” are often logical statements that directly map to a class label without relying on ML algorithms. In systems that support technical users to build a text classifier via feature-level input, features and rules often take the same form: the mentions of keywords. By incorporating features or rules into different algorithm architectures, these systems can offer varying degrees of transparency. Rules can offer the highest transparency by directly classifying text without using any ML models. Features, on the other hand, can be incorporated into transparent ML algorithms, such as Decision Trees, Naive Bayes algorithms (Kulesza et al., 2015), and Support Vector Machines (Šavelka et al., 2015). In addition, because end users might produce noisy and inconsistent features (Das et al., 2013; Boecking et al., 2020), user-generated features can also be used to label data. Weak supervision algorithms then learn from these “soft labels” (Ratner et al., 2017). However, these indirect approaches sacrifice the transparency and control that users often assume with feature-level input. For this reason, many systems designed for end users often integrate features with transparent algorithms to enhance users’ mental models of such systems (Kulesza et al., 2015; Šavelka et al., 2015). Although empirical results regarding whether feature-level input from end users improves performance have been mixed (Hema et al., 2006; Das et al., 2013; Wu et al., 2019), researchers generally agree that feature-level input requires fewer user actions to achieve comparable results to example labeling. However, features can be seen as too granular and therefore not generalizable enough, posing a challenge for end users in creating effective classifiers (Chandrasekharan et al., 2019b; Song et al., 2023).
@@ 2.2.3. Prompts
LLMs have exploded in popularity due to their proficiency in a broad range of natural language processing tasks, such as sentiment analysis and machine translation. Compared to features, natural language prompts can convey richer user guidance to models and thus promote greater generalizability across various contexts (Sahoo et al., 2024). Rather than having to train a new model for every custom task, users can simply customize LLMs by feeding them prompts at run time. Such an ability to recognize the desired task on-the-fly is called in-context learning (ICL) (Brown et al., 2020). Developing effective prompts is crucial to leveraging LLMs in content curation (Min et al., 2022; Mishra et al., 2021). To date, the most common patterns for prompting are zero-shot or few-shot prompts. Zero-shot prompts give instructions for a task without any specific examples for training (Wang et al., 2019). Research indicates that the performance of zero-shot prompts can be enhanced by iteratively refining task instructions (Efrat and Levy, 2020) or breaking down tasks into simpler subtasks (Wei et al., 2022). Alternatively, few-shot prompts incorporate a few input-output examples to showcase the desired pattern to which LLMs should adhere (Sahoo et al., 2024; Brown et al., 2020). The quality of few-shot prompts relies heavily on the selected examples (Rubin et al., 2021).
Despite the apparent simplicity of prompting, prior research has documented various challenges that impede end users in creating and refining their prompts. For example, when designing prompts for robust chatbots, users could often achieve 80% of their design goals but faced significant difficulty addressing remaining issues (Zamfirescu-Pereira et al., 2023a). Additionally, end-users often designed prompts opportunistically and thus struggled to make robust, systematic progress (Zamfirescu-Pereira et al., 2023b). Our research extends this line of work by investigating how end-users designed prompts in the setting of personal content curation. Specifically, we explore how approaches like writing rules or labeling examples can help users more easily and systematically bootstrap and iterate on their prompts.
@@ 2.3. How Existing IML Systems Overlook End User Needs in Content Curation
Researchers in End-User Programming and IML have documented various challenges end users face when building algorithms. For instance, users might struggle to find effective ways to teach the algorithm (i.e., use barrier) or diagnose why the algorithm fails (i.e., understanding barriers) (Ko et al., 2004). As a result, researchers have advocated for supporting non-experts through more intuitive approaches to model building(Yang et al., 2018). Building on this line of research, our work situates these challenges within the dynamic and evolving needs of content curation. Specifically, we argue that existing IML systems often make assumptions about end users that do not translate well into the context of content curation.